In memorium: Lajos Pukanszky (1928–1996)

I was privileged to know Lajos Pukanszky since 1976. In many ways, his personality was like his mathematics: based on a broad knowledge base, but single-minded, uncompromising, and very deep. Lajos was without a doubt the world’s foremost expert on solvable Lie groups. But his magnum opus, which occupied an entire issue of the Annales Scientifiques de l’École Normale Supérieure, was probably never read by more than a dozen people. It stands as a lofty but isolated peak on the mathematical landscape—admired by many from afar, but scaled only by a few die-hards. This was the brilliance, but also the tragedy, of Lajos Pukanszky—that his life’s work was so perfect, yet so inaccessible and underappreciated by the mathematical public.

Lajos was a truly cultured person, in a sense which we may not see again in future generations. He seemed to know all of European history, philosophy, and literature by heart. I think he identified himself with the protagonist of Thomas Mann’s Doctor Faustus, a work which he discussed with me quite a number of times. But the best quick summary of his character may be found in the quotation which he attached to the beginning of his big École Normale paper. It is a quote within a quote within a quote: Schiller, quoted by Bohr, quoted by Heisenberg. Roughly translated, part of it says “ … and in the abyss lies truth.” Pukanszky’s life work was the probing of this abyss.

Jonathan Rosenberg University of Maryland